. (a) Find the change in wavelength of 80-pm x-rays that are


scattered 120° by a target. (b) Find the angle between the directions of the recoil electron and the incident photon. (c) Find


the energy of the recoil electron.

1
Expert's answer
2022-03-01T17:32:49-0500

"\Delta \lambda=\lambda(1-\cos\varphi)=3.6~pm,"

"\theta=\arctan(1-\frac{\Delta \lambda}{\lambda})=87.4°,"

"E=hc(\frac 1{\lambda}-\frac 1{\lambda+\Delta\lambda})=9.3~\cdot 10^{-17}~J ."